# Learning with Noisy Labels: the Exploration of Error Bounds in Classification

### v1

@@ Abstract
Numerous studies have shown that label noise can lead to poor generalization performance, negatively affecting classification accuracy. Therefore, understanding the effectiveness of classifiers trained using deep neural networks in the presence of noisy labels is of considerable practical significance. In this paper, we focus on the error bounds of excess risks for classification problems with noisy labels within deep learning frameworks. We begin by exploring loss functions with noise-tolerant properties, ensuring that the empirical minimizer on noisy data aligns with that on the true data. Next, we estimate the error bounds of the excess risks, expressed as a sum of statistical error and approximation error. We estimate the statistical error on a dependent (mixing) sequence, bounding it with the help of the associated independent block sequence. For the approximation error, we first express the classifiers as the composition of the softmax function and a continuous function from $[0,1]^{d}$ to ${\mathbb{R}}^{K}$. The main task is then to estimate the approximation error for the continuous function from $[0,1]^{d}$ to ${\mathbb{R}}^{K}$. Finally, we focus on the curse of dimensionality based on the low-dimensional manifold assumption.
@@ 1 Introduction
Currently, most analyses of excess risk error bounds are focused on regression models, while research on error bounds for classification problems within deep learning frameworks remains relatively scarce, especially in the context of finite samples and label noise. This paper aims to fill this gap. We begin with noise-tolerant property of the loss functions, which demonstrates that the empirical minimizer on noisy data aligns with that on the true data. To estimate the error bound, we specifically divide them into approximation error and statistical error, which can be estimated separately. Our contributions can be summarized as follows:
(1) To the best of our knowledge, we are the first to provide an error bound for the classification problem with noisy label. The main results are in Theorems 4.1 and 6.1.
(2) The noise-tolerant property of the loss function: The noise-tolerant property guarantees the relationship of the empirical minimizers on both noisy data and true data. It is important to note that both the $\ell_{1}$ loss and the reverse cross entropy (RCE) loss exhibit noise-tolerant properties. However, for the cross entropy (CE) loss, this noise-tolerant property holds only when the classifier $f$ is a function mapping ${\mathbb{R}}^{d}$ to =$\mathcal{E}=\{e_{1},\cdots,e_{K}\}$, where $K$ represents the number of classes and $e_{i}$ is the unitary vector whose $i$-th component is one, with all other components being zero.
@@
(4) The approximation error: The classifier we aim to develop is designed to map input features to a $K$-dimensional unitary vector. To achieve this, we first express the classifiers as a composition of the softmax function and a continuous function from $[0,1]^{d}$ to ${\mathbb{R}}^{K}$. The primary objective is to estimate the approximation error for this continuous function. Although our theory emphasizes $K$-dimensional unitary outputs, it remains valid when we change the output space to $\mathcal{E}$, provided that the outputs of the true classifier and the optimal classifier are sufficiently separated from the margins.
@@
The rest of this paper is organized as follows. Section 2 reviews the related works about noisy labels and error estimations. Section 3 presents the problem setup, including the definitions of the expected excess risk estimations on the true and noisy distributions, the empirical excess risk estimations on the true and noisy data, the neural networks we considered in this paper, which is followed by the theoretical results of the excess risk under noisy labels in Section 4. Some technical details of proof are provided in Section 5. In Section 6, we present a theoretical result aimed at alleviating the curse of dimensionality. Lastly, we conclude the results in Section 7.
@@ 2.2 Error estimation
Stone [Sto82] assumed that the true regression model is $\beta$-Hölder smooth, where the smoothness index $\beta>0$. The convergence rate of the prediction error is $C_{d}n^{-2\beta/(2\beta+d)}$. This indicates that the convergence rate is exponentially dependent on the dimension $d$ of the inputs, while the prefactor $C_{d}$ is independent of the sample size $n$ but depends on $d$ and other model parameters. Research by Schmidt-Hieber [SH20] and Nakada and Imaizumi [NI20] demonstrated that deep neural network regression can achieve the optimal minimax rate established by Stone in 1982 under certain conditions. Shen, Yang and Zhang [SYZ20] derived the approximate error bound for $\beta$-Hölder continuous functions when the smoothness index $\beta\leq 1$, finding that the prefactor depends on a polynomial of $d$. Lu et al. [LSYZ21] studied the approximate error bound for $\beta$-Hölder continuous functions when the smoothness index $\beta\geq 1$, noting that its prefactor relates to an exponential form of $d$.
Jiao et al. [JSLH23] improved the result of [LSYZ21]. For the case of the smoothness index $\beta>1$, they found that the non-asymptotic upper limit of the prediction error of the empirical risk minimizer is polynomially dependent on the dimension $d$ of the predictor rather than exponentially dependent. Chen et al. [CJLZ22] focused on nonparametric regression of Hölder functions on low-dimensional manifolds using deep ReLU networks. They proved the adaptability of deep ReLU networks to low-dimensional geometric structures in data, partially explaining their effectiveness in handling high-dimensional data with low-dimensional geometric structures. Feng et al. [FJK+23] recently addressed over-parameterized deep nonparametric regression. By effectively balancing approximation and generalization errors, they derived results for Hölder functions with constrained weights, highlighting their application in reinforcement learning.
@@
3 Setup of Classification Problem with Noisy Label
We consider a $K$-class classification problem with noisy label. We start from the expected risk, which quantifies the average loss of a classifier over the entire population or distribution of data.
@@ 3 Setup of Classification Problem with Noisy Label
Let $Z=(X,Y,Y^{\eta})$ be a tuple of random vectors, where $(X,Y)$ follow a distribution $D$, $X\in{\mathbb{R}}^{d}$ represents the feature vector, $Y,Y^{\eta}\in\mathcal{Y}=\{y=[y_{1}\ \cdots\ y_{K}]^{\top}\in{\mathbb{R}}^{K}|y_{i}\in[0,1],\sum^{K}_{i=1}y_{i}=1\}$ denote the true label and the random noisy vector of $Y$, respectively. Let $f$ be a map which is from ${\mathbb{R}}^{d}$ to $\mathcal{Y}$ and $\ell(\cdot,\cdot)$ be a loss function. Then the expected risks associated with true distribution or noisy distribution are defined as
@@
A loss function, denoted $\ell(\cdot,\cdot)$, is commonly introduced to quantify the differences between $f(x)$ and its associated label $y$. This includes various loss functions such as the $\ell_{1}$ loss, the $\ell_{2}$ loss, the CE loss and the RCE loss [WMC+19], among others.
@@
as the empirical risk minimizer for the true data and the noisy data, where ${\mathcal{F}}$ represents a deep neural network class with a certain structure. Additionally, $\lfloor x\rfloor$ denotes the largest integer no greater than $x$, $\lceil x\rceil$ denotes the smallest integer no less than $x$, and $[J]=\{0,\cdots,J-1\},\ \forall J\in\mathbb{N}^{+}$.
4 The Excess Risk under Noisy Label
@@
where $X$ represents the features, $Y,Y^{\eta}$ are the true and noisy labels, respectively, $f_{0}(\cdot)$ is the unknown classifier, $n_{\eta}(\cdot)$ is the disturbing function which maps the true label to the noisy label. We assume symmetric or uniform noise. Let $\eta\in[0,1]$ denote the noise rate. For all $i,j\in\{1,\cdots,K\}$, $Y^{\eta}$ obeys the following conditional probability conditioned on $Y$
$\mathrm{Prob}(Y^{\eta}=e_{i}|Y=e_{j})=\left\{\begin{aligned} &1-\eta,&&i=j,\\ &\frac{\eta}{K-1},&&i\neq j.\end{aligned}\right.$ (4.1)
@@
We begin with an assumption regarding the logarithm function and then provide the bounds on the excess risks.
Assumption 4.1.
Let $\epsilon>0$ be a small positive number, we define
$\log x=\left\{\begin{array}[]{ll}\log x,&x\geq\epsilon,\\ A=\log\epsilon,&x\in[0,\epsilon).\end{array}\right.$
This technique in Assumption 4.1 is reasonable, which is similar to the clipping operation implemented by most deep learning frameworks.
@@
Let $\kappa$ be a continuous map from $[0,1]^{d}$ to ${\mathbb{R}}^{K}$ and define $f_{0}=\mathrm{softmax}\circ\kappa:[0,1]^{d}\to\mathcal{Y}=\{y=[y_{1}\ \cdots\ y_{K}]^{\top}|y_{i}\geq 0\ \forall i,\sum^{K}_{i=1}y_{i}=1\}$. Let ${\mathcal{F}}_{d,K}(\mathcal{W},\mathcal{D},{\mathcal{M}})$ be a class of a norm-constrained deep neural network with an input dimension $d$, an output dimension $K$, width ${\mathcal{W}}$, depth ${\mathcal{D}}$, and ${\mathcal{M}}>0$ satisfying $\|(W_{{\mathcal{D}}},b_{{\mathcal{D}}})\|_{1,\infty}\prod_{i=0}^{{\mathcal{D}}-1}\max\{\|(W_{i},b_{i})\|_{1,\infty},1\}\leq{\mathcal{M}}$, where $W_{i}\in\mathbb{R}^{p_{i+1}\times p_{i}}$ is the weight matrix, and $b_{i}\in\mathbb{R}^{p_{i+1}}$ is the bias vector in the $i$-th layer. We consider a strictly stationary $\beta$-mixing sequence defined in Definition 5.3. Let $n$ be the sample size and $a_{n},n/a_{n}$ be positive integers. For any $N,M\in\mathbb{N}^{+}$, when ${\mathcal{W}}=\max\left\{8d\left\lfloor N^{1/d}\right\rfloor+3d,16NK+30K\right\}$, ${\mathcal{D}}=11M+18$ and Assumption 4.1 holds, the excess risks for the empirical risk minimizer $\hat{f^{\eta}_{n}}\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})$ can be bounded as
(1) when the loss we consider is either the $\ell_{1}$ or RCE loss
$|\mathcal{L}(\hat{f^{\eta}_{n}})-\mathcal{L}(f_{0})|\leq$ $\frac{8\lambda K{\mathcal{M}}\sqrt{{\mathcal{D}}+2+\log d}}{\sqrt{na_{n}}}+\frac{4\lambda Kn\beta_{a_{n}}}{a_{n}}$
$\qquad+131c\lambda\sqrt{d}\left\|w_{\kappa}\left((N^{2}M^{2}\log_{3}(N+2))^{-1/d}\right)\right\|_{1},$ (4.2)
(2) when the loss function we consider is one of the $\ell_{1}$, CE and RCE losses
$|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}(f_{0})|\leq$ $\frac{8\lambda K{\mathcal{M}}\sqrt{{\mathcal{D}}+2+\log d}}{\sqrt{na_{n}}}+\frac{4\lambda Kn\beta_{a_{n}}}{a_{n}}$
$\qquad+131c\lambda\sqrt{d}\left\|w_{\kappa}\left((N^{2}M^{2}\log_{3}(N+2))^{-1/d}\right)\right\|_{1},$ (4.3)
where $w_{f}(s)=\left\{\sup\left|f(x)-f(y)\right|:x,y\in[0,1]^{d},\left\|x-y\right\|_{2}\leq s\right\}$ for any continuous function $f$ and $w_{\kappa}(s)=[w_{\kappa^{(1)}}(s)\ \cdots\ w_{\kappa^{(K)}}(s)]^{\top}$, $\beta_{a_{n}}$ is the $a_{n}$-th $\beta$-mixing coefficient of samples defined in Definition 5.3, $c$ is a positive constant and
$\lambda=\left\{\begin{array}[]{ll}1,&\ell_{1}\text{\ loss},\\ \frac{1}{\epsilon},&\text{CE loss},\\ -A,&\text{RCE loss}.\end{array}\right.$
@@
In general, both the optimal classifier and the learned classifier within the framework of deep neural networks aim to maximize separability between different classes while minimizing in-class distances, which enables the classifier to achieve maximum margin classification. Furthermore, in $K$-class classification tasks, the theory of neural collapse for the CE loss [ZDZ+21, TKVB22, Liu24] suggests that the last-layer classifiers of well-trained deep neural networks, when trained on a balanced dataset, correspond to the $K$ vertices of the Simplex Equiangular Tight Frame (ETF). This framework ensures both concentration within classes and separation between different classes. Although we do not concentrate on $f_{0}:[0,1]^{d}\to\mathcal{E}=\{e_{1},\cdots,e_{K}\}$ in relation to the CE loss, the results on the empirical risk for the true distribution hold true by Theorem 4.1, provided that the outputs of the true classifier and the optimal classifier are sufficiently separated from the margins.
@@
Equation (4.2) in Theorem 4.1 is based on the noise-tolerant property of the loss function, which guarantees $\hat{f^{\eta}_{n}}=\hat{f_{n}}$. That is, $\mathcal{L}(\hat{f^{\eta}_{n}})-\mathcal{L}(f_{0})=\mathcal{L}(\hat{f_{n}})-\mathcal{L}(f_{0})$. It is important to note that both the $\ell_{1}$ loss and the RCE loss exhibit noise-tolerant properties. However, for the CE loss, this noise-tolerant property holds only when $f$ is a function mapping ${\mathbb{R}}^{d}$ to =$\mathcal{E}=\{e_{1},\cdots,e_{K}\}$. We will discuss this further in Subsection 5.1.
Additionally, Assumption 4.1 is required only for the CE and RCE losses; it is not needed for the $\ell_{1}$ loss at all.
Furthermore, we have
$w_{f}\left((N^{2}M^{2}\log_{3}(N+2))^{-1/d}\right)\leq O((N^{2}M^{2}\log_{3}(N+2))^{-1/d})$
for any uniformly continuous functions $f(x)$.
Lastly, $\beta_{a_{n}}$ tends to zero as $a_{n}$ approaches infinity. Thus, $\beta_{a_{n}}$ can be made small by choosing $a_{n}$ sufficiently large.
@@
In this subsection, we provide a detailed proof of Theorem 4.1. In Subsection 5.1, we discuss the noise-tolerant properties of the loss function, demonstrating that the empirical minimizer on noisy data aligns with that on the true data. In Subsection 5.2, we estimate the error bounds of the excess risk for different loss functions, categorizing them into statistic error and approximation error. We analyze these two types of errors in Subsections 5.3 and 5.4.
5.1 Noise-tolerance
In the following, we will demonstrate that the empirical risk minimizer for the noisy data is the same as that for the true data, which is referred as noise-tolerant.
Definition 5.1.
To prove the noise-tolerant property, it is essential to define the symmetric loss function. We claim that a loss function is noise-tolerant if it is a symmetric loss function, which is stated in Theorem 5.1.
Definition 5.2 (Equation (2) in [GKS17]).
Theorem 5.1.
Let $K$ be the number of classes in the classification problem with noisy labels. We assume that the noisy labels are derived from the true labels through symmetric noise, as defined in (4.1), with the noise rate $\eta<\frac{K-1}{K}$. When a loss function satisfies (5.1), it is noise-tolerant.

### v2

# The Exploration of Error Bounds in Classification with Noisy Labels

@@ Abstract
Numerous studies have shown that label noise can lead to poor generalization performance, negatively affecting classification accuracy. Therefore, understanding the effectiveness of classifiers trained using deep neural networks in the presence of noisy labels is of considerable practical significance. In this paper, we focus on the error bounds of excess risks for classification problems with noisy labels within deep learning frameworks. We derive error bounds for the excess risk, decomposing it into statistical error and approximation error. To handle statistical dependencies (e.g., mixing sequences), we employ an independent block construction to bound the error, leveraging techniques for dependent processes. For the approximation error, we establish these theoretical results to the vector-valued setting, where the output space consists of $K$-dimensional unit vectors. Finally, under the low-dimensional manifold hypothesis, we further refine the approximation error to mitigate the impact of high-dimensional input spaces.
@@ 1 Introduction
Currently, research on error bounds for classification problems within deep learning frameworks remains relatively scarce, especially in the context of finite samples and label noise. This paper aims to fill this gap. To estimate the error bound, we specifically divide them into approximation error and statistical error, which can be estimated separately. Our contributions can be summarized as follows:
(1) To the best of our knowledge, we are the first to provide an error bound for the classification problem with noisy labels. The main results are in Theorem 4.1, Corollary 4.3 and Theorem 7.1.
@@
(3) The approximation error: While prior works primarily focused on scalar-valued functions (output space ${\mathbb{R}}^{1}$), we generalize these theoretical foundations to the vector-valued setting, where the output space consists of $K$-dimensional unit vectors.
@@
The rest of this paper is organized as follows. Section 2 reviews the related works about noisy labels and error estimations. Section 3 presents the problem setup, including the definitions of the expected excess risk estimations on the true and noisy distributions, the empirical excess risk estimations on the true and noisy data, the neural networks we considered in this paper, which is followed by the theoretical results of the excess risk under noisy labels in Section 4. Some technical details of proofs are provided in Sections 5 and 6. In Section 7, we present a theoretical result aimed at alleviating the curse of dimensionality. Lastly, we conclude the results in Section 8.
@@ 2.2 Error estimation
Stone [Sto82] assumed that the true regression model is $\beta$-Hölder smooth, where the smoothness index $\beta>0$. The convergence rate of the prediction error is $C_{d}n^{-2\beta/(2\beta+d)}$. This indicates that the convergence rate is exponentially dependent on the dimension $d$ of the inputs, while the prefactor $C_{d}$ is independent of the sample size $n$ but depends on $d$ and other model parameters. Research by Schmidt-Hieber [SH20] and Nakada and Imaizumi [NI20] demonstrated that deep neural network regression can achieve the optimal minimax rate established by Stone in 1982 under certain conditions. Shen, Yang and Zhang [SYZ20] derived the approximate error bound for $\beta$-Hölder continuous functions when the smoothness index $\beta\leq 1$, finding that the prefactor depends on a polynomial of $d$. Lu et al. [LSYZ21] studied the approximate error bound for $\beta$-Hölder continuous functions when the smoothness index $\beta\geq 1$, noting that its prefactor relates to an exponential form of $d$. Jiao et al. [JSLH23] improved the result of [LSYZ21]. For the case of the smoothness index $\beta>1$, they found that the non-asymptotic upper limit of the prediction error of the empirical risk minimizer is polynomially dependent on the dimension $d$ of the predictor rather than exponentially dependent. Chen et al. [CJLZ22] focused on nonparametric regression of Hölder functions on low-dimensional manifolds using deep ReLU networks. They proved the adaptability of deep ReLU networks to low-dimensional geometric structures in data, partially explaining their effectiveness in handling high-dimensional data with low-dimensional geometric structures. Feng et al. [FJK+23] recently addressed over-parameterized deep nonparametric regression. By effectively balancing approximation and generalization errors, they derived results for Hölder functions with constrained weights, highlighting their application in reinforcement learning.
@@
3 Setup of Classification Problem with Noisy Labels
We consider a $K$-class classification problem with noisy labels. We start from the expected risk, which quantifies the average loss of a classifier over the entire population or distribution.
@@ 3 Setup of Classification Problem with Noisy Labels
Let $Z=(X,Y,Y^{\eta})$ be a tuple of random vectors, where $(X,Y)$ follow a distribution $D$, $X\in{\mathbb{R}}^{d}$ represents the feature vector, $Y,Y^{\eta}\in\mathcal{Y}=\{y=[y_{1}\ \cdots\ y_{K}]^{\top}\in{\mathbb{R}}^{K}|y_{i}\in[0,1],\sum^{K}_{i=1}y_{i}=1\}$ denote the true label and the perturbed noisy vector of $Y$, respectively. Let $f$ be a map which is from ${\mathbb{R}}^{d}$ to $\mathcal{Y}$ and $\ell(\cdot,\cdot)$ be a loss function. Then the expected risks associated with true distribution or noisy distribution are defined as
@@
A loss function, denoted $\ell(\cdot,\cdot)$, is commonly introduced to quantify the differences between $f(x)$ and its associated label $y$. This includes various loss functions such as the $\ell_{1}$ loss, the $\ell_{2}$ loss, the cross entropy (CE) loss and the reverse CE loss [WMC+19], among others.
@@
as the empirical risk minimizer for the true data and the noisy data, where ${\mathcal{F}}$ represents a deep neural network class with a certain structure. Additionally, $\lfloor x\rfloor$ denotes the largest integer no greater than $x$, $\lceil x\rceil$ denotes the smallest integer no less than $x$, $x\lesssim y$ or $y\gtrsim x$ denotes $x\leq Cy$ for some positive constant $C>0$ and $x\asymp y$ means $x\lesssim y\lesssim x$. For a multi-index $s=[s_{1}\cdots s_{d}]$ and a $d$-dimensional vector $x=[x_{1}\cdots x_{d}]\in{\mathbb{R}}^{d}$, $x^{s}=x^{s_{1}}_{1}\cdots x^{s_{d}}_{d}$. $\mathbb{N}$ and $\mathbb{N}^{+}$ are the non-negative and positive integer spaces, respectively.
4 The Excess Risk under Noisy Labels
@@
where $X$ represents the features, $Y,Y^{\eta}$ are the true and noisy labels, respectively, $f_{0}(\cdot)$ is the unknown classifier, $n_{\eta}(\cdot)$ is the disturbing function which maps the true labels to the noisy labels.
@@
For the unknown classifiers, we represent as $f_{0}=\mathrm{softmax}\circ\kappa_{0}:[0,1]^{d}\to\mathcal{Y}=\{y=[y_{1}\ \cdots\ y_{K}]^{\top}|y_{i}\geq 0\ \forall i,\sum^{K}_{i=1}y_{i}=1\}$, where $\kappa_{0}$ belongs to the following class of smooth maps:
Definition 4.1.
We define the following class of smooth maps:
$\mathcal{C}_{\tau,d,K,B}:=\left\{f=[f_{1}\cdots f_{K}]^{\top}:f_{i}\in C^{r}\left([0,1]^{d}\right):\left\|f_{i}\right\|_{C^{0,\tau}}\leq B,\forall i\right\},$
where $\tau\in(0,\infty)$ with $\tau=r+\nu$, $r\in\mathbb{N}$, $d\in\mathbb{N}^{+}$, $\nu\in(0,1]$ and the $C^{0,\tau}$-norm is defined by
$\left\|g\right\|_{C^{0,\tau}}:=\max\left\{\max_{\left|\alpha\right|\leq r}{\sup_{x\in[0,1]^{d}}\left|\partial^{\alpha}g\right|},\max_{\left|\alpha\right|=r}\sup_{x,y\in[0,1]^{d},x\neq y}\frac{\partial^{\alpha}g(x)-\partial^{\alpha}g(y)}{\left|x-y\right|^{\nu}}\right\},\ \forall g\in C^{r}\left([0,1]^{d}\right).$
In practice, the samples in a dataset may not be independent. In this paper, we focus on a dependent (mixing) sequence.
Definition 4.2 ($\beta$-mixing, Subsection 2.2 in [Yu94]).
@@
Let $f_{0}=\mathrm{softmax}\circ\kappa_{0}:[0,1]^{d}\to\mathcal{Y}=\{y=[y_{1}\ \cdots\ y_{K}]^{\top}|y_{i}\geq 0\ \forall i,\sum^{K}_{i=1}y_{i}=1\}$ and $\kappa_{0}\in\mathcal{C}_{\tau,d,K,B}$. Let $S^{\eta}=\left\{z^{\eta}_{i}=(x_{i},y^{\eta}_{i})\right\}_{i=1}^{n}$ be a strictly stationary $\beta$-mixing sequence and $(a_{n},\mu_{n})$ be an integer pair such that $n=a_{n}\mu_{n}$. For a large enough ${\mathcal{M}}\in\mathbb{N}^{+}$, when ${\mathcal{W}}\gtrsim{\mathcal{M}}^{d/(d+1)}\log{\mathcal{M}}$, ${\mathcal{D}}\gtrsim\log{\mathcal{M}}$ and $\left|\ell(\mathrm{softmax}(a),q)-\ell(\mathrm{softmax}(b),q)\right|\leq\lambda\|a-b\|_{2},\ \forall a,b\in{\mathbb{R}}^{K},\forall q\in\mathcal{Y}$, the excess risks for the empirical risk minimizer $\hat{f^{\eta}_{n}}\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})$ can be bounded as
$\mathbb{E}_{S^{\eta}}[|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}(f_{0})|]\lesssim$ $\frac{8\lambda\sqrt{K}{\mathcal{M}}\sqrt{{\mathcal{D}}+2+\log d}}{\sqrt{na_{n}}}+\frac{4\lambda\sqrt{K}n\beta_{a_{n}}}{a_{n}}+\lambda\sqrt{K}{\mathcal{M}}^{-\tau/(d+1)},$ (4.2)
where $\beta_{a_{n}}$ is the $a_{n}$-th $\beta$-mixing coefficient of samples defined in Definition 4.2.
@@
The condition $\left|\ell(\mathrm{softmax}(a),q)-\ell(\mathrm{softmax}(b),q)\right|\leq\lambda\|a-b\|_{2},\ \forall a,b\in{\mathbb{R}}^{K},\forall q\in\mathcal{Y}$ holds for all the $\ell_{p},p\geq 1$ losses, the CE loss and the reverse CE loss, see Appendix B.
When the loss function we consider is noise-tolerant, we have a more strong theoretical result.
Definition 4.3.
Corollary 4.3.
Under the same assumption on Theorem 4.1, when the loss we consider is noise-tolerant, we have
$\mathbb{E}_{S}[|\mathcal{L}(\hat{f^{\eta}_{n}})-\mathcal{L}(f_{0})|]\lesssim$ $\frac{8\lambda\sqrt{K}{\mathcal{M}}\sqrt{{\mathcal{D}}+2+\log d}}{\sqrt{na_{n}}}+\frac{4\lambda\sqrt{K}n\beta_{a_{n}}}{a_{n}}+\lambda\sqrt{K}{\mathcal{M}}^{-\tau/(d+1)},$ (4.3)
where $\beta_{a_{n}}$ is the $a_{n}$-th $\beta$-mixing coefficient of samples defined in Definition 4.2.
Proof.
We defer the proof to Section 6. ∎
Remark 4.4.
Equation (4.3) in Corollary 4.3 is based on the noise-tolerant property of the loss function, which guarantees $\hat{f^{\eta}_{n}}=\hat{f_{n}}$. That is, $\mathcal{L}(\hat{f^{\eta}_{n}})-\mathcal{L}(f_{0})=\mathcal{L}(\hat{f_{n}})-\mathcal{L}(f_{0})$.
@@
In this subsection, we provide a detailed proof of Theorem 4.1. In Subsection 5.1, we estimate the error bounds of the excess risk for different loss functions, categorizing them into statistic error and approximation error. We analyze these two types of errors in Subsections 5.2 and 5.3, which are followed by the proof of Theorem 4.1.
5.1 Error analysis
In this subsection, we analyze the error bounds for excess risks on noisy and true data.
Lemma 5.1.
Let $f_{0}=\mathrm{softmax}\circ\kappa_{0}$ represent the true classifier, and $\hat{f^{\eta}_{n}}$ denote the empirical risk minimizer on the noisy data. Let $|\ell(\mathrm{softmax}(a),q)-\ell(\mathrm{softmax}(b),q)|\leq\lambda\|a-b\|_{2},\ \forall a,b\in{\mathbb{R}}^{K},\forall q\in\mathcal{Y}$. We have the following
$|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}(f_{0})|\leq 2\sup_{\begin{subarray}{c}f=\mathrm{softmax}\circ\kappa,\\ \kappa\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D},{\mathcal{M}})\end{subarray}}\left|\mathcal{L}^{\eta}(f)-\mathcal{L}^{\eta}_{n}(f)\right|+\lambda\inf_{\kappa\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D},{\mathcal{M}})}\left\|\kappa-\kappa_{0}\right\|_{L^{2}([0,1]^{d})}.$ (5.1)